OpenStudy (anonymous):

The surface area of a beach ball is 144π in2. Find the volume of air the beach ball will hold when it is fully inflated. A) 16π in3 B) 36π in3 C) 144π in3 D) 288π in3

pooja195 (pooja195):

Surface Area of a sphere = 4πr² = 144π in²

pooja195 (pooja195):

refresh your page

pooja195 (pooja195):

4πr² = 144π r² = 144π/4π = 36 r = √36 = 6 in

pooja195 (pooja195):

volume of the beach ball (sphere) is = (4/3)πr³ Volume of the air the beach ball will hold when fully inflated = (4/3)πr³ = (4/3)π(6)³ = 288π in³

OpenStudy (anonymous):

thanks